OpenStudy (anonymous):

Let f(x)=sqrt{x-5} and g(x)=x-3. Find (f/g) (9).

OpenStudy (anonymous):

possibilities: (a) i/6 (b) 1/3 (c) 2 (d) 0 (e) 1

OpenStudy (anonymous):

sqrt (9-5) = 4 9-3 = 6 4/6 = 1/3 b

OpenStudy (anonymous):

you sure u did this right? ...4/6 reduced is 2/3 not 1/3

OpenStudy (anonymous):

oh sorry sqrt(9-5) = 2 so 2/6 = 1/3

OpenStudy (anonymous):

ok thanks